OpenStudy (anonymous):

y=d/dx integralcosx to 0 t/(1+t)dt

OpenStudy (anonymous):

\[d/dx \int\limits_{0}^{cosx} \frac{ t }{ 1+t } dt\]

OpenStudy (anonymous):

\[\frac{cosx}{1+cosx}\]

OpenStudy (anonymous):

how did u do it?

OpenStudy (anonymous):

@Outkast3r09 I think u made a miss take in your first step (While u made the integral)

OpenStudy (anonymous):

the derivative of an integral is simply the integral itself with the variables in other words ... integrals and derivatives cancel eachother out

OpenStudy (anonymous):

\[\frac{cosx}{1+cosx} \]is your answer

OpenStudy (anonymous):

It should be\[\cos(x)-\ln(1+\cos(x))\]

OpenStudy (anonymous):